solve the following initial value problem d^2y/dx^2 + dy/dx -2y=-6 sin2x-18cos2x where y(0)=2, y'(0)=2
1
Expert's answer
2020-05-06T19:09:09-0400

The given differential equation is

"\frac{d^2y}{dx^2}+\frac{dy}{dx}-2y=-6 \ sin2x-18 \ cos2x" .


The auxiliary equation is "m^2+m-2=0"

"\implies m^2+2m-m-2=0"

"\implies m(m+2)-1(m+2)=0"

"\implies (m+2)(m-1)=0"

"\implies m=1,-2"

"\therefore C.F.=A\ e^x+B\ e^{-2x}"


Now , "P.I.=\frac{1}{D^2+D-2}(-6\ sin2x-18\ cos2x)"

"=\frac{1}{D^2+D-2}(-6\ sin2x)+\frac{1}{D^2+D-2}(-18\ cos2x)"


"= \frac{-6}{-4+D-2}(sin2x)+\frac{-18}{-4+D-2}(cos2x)"


"=\frac{-6}{D-6}(sin2x)-\frac{18}{D-6}(cos2x)"


"=-6×\frac{D+6}{D^2-36}(sin2x)-18×\frac{D+6}{D^2-36}(cos2x)"


"=\frac{6}{40}×(2\ cos2x+6\ sin2x)+\frac{18}{40}×(-2\ sin2x +6\ cos2x)"


"=(\frac{18}{20}+\frac{-18}{20})sin2x+(\frac{6}{20}+\frac{54}{20})cos2x"


"=3 \ cos2x" .

Therefore the complete solution is

"y=A\ e^x+B \ e^{-2x}+3\ cos2x" ...................(1)

Differentiating (1) ,we get

"y'=A\ e^x-2B\ e^{-2x}-6\ sin2x" .............(2)


But given that "y(0)=2" and "y'(0)=2" .

Therefore from equation (1) and (2) ,we get

"A+B+3=2 \ ,\implies A+B=-1" .............(3)


"A-2B=2" .........(4)

Subtracting (4) from (3) ,we get

"3B=-3 \ , \implies B=-1"

Putting the value of B in equation (3) ,we get

A=0.

"\therefore y=-e^{-2x}+3 \ cos2x"